. The acceleration of a particle P moving in a straight line is (t

2 − 9t + 18)ms−2

,

where t is the time in seconds.

(a) Find the values of t for which the acceleration is zero.

(b) It is given that when t = 3 the velocity of P is 9ms−1

.Find the velocity of P when t =0

(c) Show that the direction of motion of P changes before t = 1.


1
Expert's answer
2021-10-31T18:16:02-0400

a)

"t^2-9t+18=0,"

"t=3 ~\text{or} ~t=6,"

b)

"v(t)=\int adt=\frac 13t^3-\frac 92t^2+18t+C,"

"v(3)=23.5+C=9,\implies" "C=-14.5,"

"v(0)=-14.5~\frac ms,"

c)

"\frac 13t^3-\frac 92t^2+18t-14.5=0,"

"t\approx 1.07>1."